OpenStudy (anonymous):

Which of the following is the radical expression of a5/7 ?

OpenStudy (anonymous):

when you have a fractional exponent the denominator is the index or root and the numerator is the exponent; the exponent can go inside or outside the radical:\[a^{5/7}=\sqrt[7]{a^5}=\left( \sqrt[7]{a} \right)^5\]

jhonyy9 (jhonyy9):

\[\sqrt[7]a ^{5}\]